# The Orientation Boundary for Step-Duplicating Recursors: Mechanized Impossibility, Escape, and Certification

### 19 Mar 2026 v7

@@ 4.7 Proof-theoretic context
Machine-checked in Meta/StepDuplicatingSchema.lean.
This is the precise sense in which path-order success lies outside the present Nat-valued measure framework: not because path orders are unsound, but because successful orienters cannot be represented by the formal classes isolated above.
@@
Machine-checked in Meta/PolyInterpretation_FullStep.lean as W_orients_step, wf_StepRev_poly, W_violates_transparency, W_not_additive, and W_not_affine.
@@
For the fixed KO7 signature, precedence, and same-head clause formalized in Meta/MPO_FullStep.lean, the reverse root relation induced by Step is well-founded. Machine-checked as wf_MPORev together with the induced reverse-Step theorem wf_StepRev_mpo. This is a fixed-signature KO7 result, not a generic MPO well-foundedness theorem over arbitrary signatures.
@@
The fixed-signature KO7 MPO proof is witnessed by an ordinal ranking into the initial segment below $\varphi_{7}(0)$, that is, below Ordinal.veblen70. This gives a concrete proof-theoretic upper bound for the specialized MPO argument without appealing to generic path-order metatheory over arbitrary signatures. Machine-checked in Meta/MPO_ProofTheoreticBound.lean as mpoOrd_lt_mpoVeblenBound and wf_StepRev_mpo_below_veblen7.
@@
Consequently that bad precedence cannot orient all root steps of Step. Machine-checked in Meta/MPO_Precedence_Barrier.lean as not_mpoBad_rec_succ_instance and no_global_step_orientation_mpo_bad_prec.
@@
Theorem universe. The following theorem is stated over an explicit direct universe consisting of twelve listed families: additive compositional, transparent-compositional, pumped affine, pumped restricted-quadratic, pumped bounded cross-term quadratic, pumped bounded multilinear, pumped generalized bounded-polynomial, pumped max-plus, tracked-primary componentwise pair, tracked-primary lexicographic pair, KO7-specific max-depth, and KO7-specific pure head-precedence. For the pair-valued families, the theorem reads their first component as the tracked primary scalar. The Lean development first proves the scalar-only preliminary ko7_nat_direct_escape_trichotomy; the theorem stated here is the KO7-specific extension ko7_direct_escape_trichotomy_extended, which adds the tracked-primary pair families. A second theorem, ko7_projection_escape_trichotomy, extends the same classification pattern to successful componentwise orienters that admit a fixed scalar projection whose strict decrease is forced by the ambient order and whose projected scalar lies in the listed Nat-valued universe. This is how the weighted scalar-projection and balanced mixed-coordinate matrix families are absorbed into the trichotomy layer. The generalized degree-bounded polynomial family of Theorem 4.6 is now internalized directly into the Nat-valued universe; projection-based matrix families still enter through the separate projection theorem. Dependency-pair frameworks, generic path-order metatheory, arbitrary semantic methods, and arbitrary mixed matrix orders remain outside this theorem universe.
@@
Equivalently, every successful direct root-step orienter in the theorem universe escapes by violating wrapper sensitivity on its tracked primary scalar, violating transparency there, or leaving the formalized direct families. Machine-checked in Meta/EscapeTrichotomy.lean as ko7_direct_escape_trichotomy_extended; the scalar-only preliminary theorem is ko7_nat_direct_escape_trichotomy.
Projection-based extension. The same escape pattern also covers the weighted scalar-projection componentwise family and the balanced mixed-coordinate componentwise family. The reason is structural rather than ad hoc: if the ambient componentwise order forces strict decrease of a fixed scalar projection, and that projected scalar already belongs to the explicit Nat-valued direct universe above, then the trichotomy applies to that scalar as well. Machine-checked in Meta/EscapeTrichotomy.lean as ko7_projection_escape_trichotomy, with the concrete representability lemmas matrixFunctional_projection_representable and matrixMix2_sum_projection_representable.
Scope note. The theorem is intentionally restricted to a finite, reviewable direct universe plus a separate projection-based extension principle. This is a deliberate design choice. A theorem over a listed class is verifiable and falsifiable, while an unbounded claim about all imaginable orienters would be either tautological, by defining “outside the formalized classes” to absorb every counterexample, or unprovable with current methods. The bounded cross-term, bounded multilinear, generalized bounded-polynomial, and max-plus scalar families are included directly; the weighted scalar-projection and balanced mixed-coordinate matrix families are included through the projection theorem; and the current matrix-side strengthening now also exposes fixed-row and row-sum projection corollaries in Meta/MatrixProjectionCoverage.lean. Dependency-pair frameworks, generic path-order metatheory, arbitrary semantic methods, and arbitrary mixed matrix orders remain outside the present theorem universe, even though the paper separately proves negative or positive results for some of them.
@@ 4.8 Scope
Status of full Step
Root-step Step is terminating: mechanized via the nonlinear polynomial interpretation (wf_StepRev_poly) and the KO7-specialized MPO (wf_StepRev_mpo). The full context closure is also strongly normalizing in Lean via wf_StepCtxFullRev_poly, and independently via TTT2/CeTA (§6). Full Step is not locally confluent.
@@
For every trace $a$, the full kernel relation fails local join at $\mathrm{eqW}\;a\;a$: the two one-step reducts $\mathrm{void}$ and $\mathrm{integrate}(\mathrm{merge}\;a\;a)$ are distinct full-step normal forms.
Machine-checked in Meta/EqW_Guard_Barrier.lean as not_localJoinStep_eqW_refl. This is why the SafeStep eqW guards are genuine confluence guards rather than presentation artifacts.
@@
The barrier section isolates a concrete frontier. Inside the formalized direct classes, the duplicating step is excluded by the schema, affine, restricted quadratic, bounded cross-term quadratic, bounded multilinear, generalized bounded polynomial, max-plus, tracked componentwise fixed-dimension vector, tracked lexicographic pair, balanced mixed-coordinate pair, and weighted scalar-projection componentwise barriers, together with the symbolic variable-condition / KBO corollary. The trichotomy theorem classifies the scalar side through max-plus and multilinear, and extends to the weighted-functional and balanced mixed-coordinate matrix families through projected representability, while the generalized polynomial theorem isolates the remaining bounded direct-polynomial escape condition. Crossing the frontier therefore requires one of three moves exhibited in the artifact: projection-style transformation (DP), direct nonlinear escape outside the formalized classes (the global witness $W$ and the KO7-specialized MPO), or ablation of the offending structure itself (removing duplication dissolves the barrier, while a directed wrapper-collapse surrogate does not). The SCC theorems show that this frontier is not confined to immediate single-rule duplication. Delayed duplication across one alternating cycle already recreates the additive and affine obstruction, and even a multiplicity-preserving SCC can force additive failure once a synchronized cycle exposes the latent payload under visible wrappers.
@@ 4.9 Orientation catalog
    Path Orders ($\dagger$/L): LPO orients the full TRS (CeTA-certified); this is a framework escape from the $\mathbb{N}$-valued measure classes (Cor. 4.25). Meta/MPO_FullStep.lean proves KO7-specialized MPO orientation (wf_MPORev, wf_StepRev_mpo); this is a fixed-signature comparator, not a generic MPO metatheorem.
@@ 4.10 Gallery of explicitly excluded methods
The Ladder Paradox (exact polynomial tie).
@@
Machine-checked as poly_mul_ties_rec_succ. This is not merely non-strict failure but exact cancellation: the duplication of $s$ under app is algebraically absorbed by the multiplicative coupling. Only nonlinear cross-variable coupling (as in $W$) breaks this tie.
@@
The duplication barrier and the flag barrier target different rules and impose contradictory demands on any single $\mathbb{N}$-valued measure: rec_succ requires insensitivity to duplication of $s$ (additive size fails), while merge_void requires sensitivity to flag structure (the $\delta$-flag must strictly decrease). Machine-checked as dual_barrier_rec_succ_and_merge_void.
@@
In a strongly normalizing and confluent calculus with a computable normalizer and decidable equality, the set $\{t\mid t\Rightarrow^{*}c\}$ for any normal-form target $c$ is decidable: compute $\mathrm{normalizeSafe}(t)$ and compare. This is formalized as reachability_decidable in Meta/Newman_Safe.lean: given $c$ in safe normal form, the predicate $t\mapsto\texttt{SafeStepStar}\;t\;c$ is a DecidablePred, reducing reachability to a single normalization followed by equality comparison. Complexity bounds for related TRS classes are discussed in [2, 6, 15].
@@ 5.1 Strong normalization (SafeStep)
Proof. By rule head. For rec-succ, $\delta\text{-flag}$ drops ($1\to 0$). For rec-zero, $\kappa^{M}$ drops (DM). For tie branches, $\tau$ resolves.
@@
with additive or constant clauses for the remaining constructors. The constant +5+5 in the exponent is load-bearing: the central lemma ctxFuel_rec_succ_drop shows that the duplicating step $\texttt{rec}\Delta(b,s,\delta(n))\to\texttt{app}(s,\texttt{rec}\Delta(b,s,n))$ strictly decreases ctxFuel, because $\texttt{ctxFuel}(s)+1<2^{\texttt{ctxFuel}(s)+5}$.
@@
For the partial context closure SafeStepCtx, the remaining global local-join obligation is discharged, so SafeStepCtx is confluent. The exact Newman characterization is still available:
@@
Since strong normalization is already proved, this gives both an unconditional contextual confluence theorem and the exact Newman-style equivalence for the partial context closure.
Machine-checked in Meta/SafeStepCtx_Confluence.lean as localJoinAll_ctx, confluentSafeCtx, confluentSafeCtx_iff_localJoinAll, together with the specialized Newman direction newman_safeCtx.
@@
Machine-checked in Meta/ContextClosed_SN_Full.lean as wf_StepCtxFullRev_poly. The proof lifts the nonlinear polynomial witness $W$ through every constructor position and shows that every contextual contraction still strictly decreases $W$.
@@
Machine-checked in Meta/SafeStep_Complexity.lean as stepCtxFullPow_length_le_W and stepCtxFullPow_length_lt_W. This is a concrete complexity theorem for the full contextual relation; the ordinal discussion below remains a separate proof-theoretic calibration for the guarded SafeStep argument.
@@ 5.2 Certified normalizer
Note that normalizeSafe computes normal forms with respect to the guarded relation SafeStep, not the full unguarded Step; a SafeStep-normal form may still admit unguarded reductions. Moreover, SafeStep is deterministic at the root (safeStep_deterministic in Meta/NormalizeSafe_LowerBound.lean), so the certified normalizer follows the unique possible guarded root reduction at each step. Under confluence, reachability reduces to normalization and equality check. For safe normal-form targets, the resulting decision procedure also carries an explicit certified cost envelope: Meta/Reachability_Complexity.lean proves reachabilityDecisionCost_le_complexity_bound, together with the linear lower-bound family reachabilityDecision_has_linear_lower_family.
@@
Machine-checked in Meta/NormalizeSafe_LowerBound.lean via mergeVoidChain_pow_void, normalizeSafe_mergeVoidChain, and normalizeSafeSteps_mergeVoidChain. This gives a concrete linear lower bound for the cost of the certified root normalizer on one explicit family.
@@
For every term $t$, the deterministic root-step counter normalizeSafeStepstt is realized by an exact guarded root reduction path and is bounded above by the same explicit size-based tower bound used for SafeStepCtx:
$\texttt{normalizeSafeSteps}(t)\;\leq\;\mathrm{complexity\_bound}(\mathrm{termSize}(t)).$
Machine-checked in Meta/SafeRoot_Complexity.lean; the exact theorem identifiers are listed in the claim-to-code appendix. This yields a fully mechanized algorithmic envelope for the certified root normalizer: an exact linear lower-bound family on merge spines and a computable upper bound inherited from the context-closed guarded measure.
In particular, the guarded root relation has a concrete linear lower-bound family for derivation length: each step from $m_{n+1}$ to $m_{n}$ removes exactly one outer mergevoid layer, so the family grows by one constructor layer per step while requiring exactly $n$ guarded root contractions. The current upper-envelope propositions are still coarse: they route the root normalizer and the partial-context closure through ctxFuel and explicit size-indexed envelopes. The remaining open problem is therefore no longer the absence of any upper theorem, but tightening those explicit envelopes toward a sharper slow-growing extraction from the ordinal calibration below.

### 08 Apr 2026 v8

@@ 4.7 Proof-theoretic context
See Meta/StepDuplicatingSchema.lean.
Here and below, the schema-level API uses the prefix no_global_orients_*, while the KO7-instantiation API uses no_global_step_orientation_*. This is the precise sense in which path-order success lies outside the present Nat-valued measure framework: not because path orders are unsound, but because successful orienters cannot be represented by the formal classes isolated above.
@@
The Lean proof is in Meta/PolyInterpretation_FullStep.lean (W_orients_step, wf_StepRev_poly, W_violates_transparency, W_not_additive, W_not_affine).
In that interpretation the eqW clause is $W(\mathrm{eqW}\;a\;b)=W(a)+W(b)+3$. The constant +3+3 is tight: R_eq_diff needs $W(\mathrm{eqW}\;a\;b)>W(\mathrm{integrate}(\mathrm{merge}\;a\;b))=W(a)+W(b)+2$, while R_eq_refl needs $W(\mathrm{eqW}\;\mathrm{void}\;\mathrm{void})>W(\mathrm{void})=1$. So $3$ is the smallest uniform choice that makes both rules strictly decreasing.
@@
[…]
@@
The fixed-signature KO7 MPO proof is witnessed by an ordinal ranking into the initial segment below $\varphi_{7}(0)$, that is, below Ordinal.veblen70. Here $\varphi_{\alpha}$ denotes the Veblen hierarchy, so $\varphi_{7}(0)$ lies below the Feferman–Schütte ordinal $\Gamma_{0}$. This gives a concrete proof-theoretic upper bound for the specialized MPO argument without appealing to generic path-order metatheory over arbitrary signatures. Proved as mpoOrd_lt_mpoVeblenBound and wf_StepRev_mpo_below_veblen7 in Meta/MPO_ProofTheoreticBound.lean.
@@
Consequently that bad precedence cannot orient all root steps of Step. See Meta/MPO_Precedence_Barrier.lean (not_mpoBad_rec_succ_instance, no_global_step_orientation_mpo_bad_prec).
@@
Theorem universe. The following theorem is stated over an explicit direct universe consisting of twelve listed families, namely the twelve-family trichotomy universe distinct from the twelve-family barrier package discussed above: additive compositional, transparent-compositional, pumped affine, pumped restricted-quadratic, pumped bounded cross-term quadratic, pumped bounded multilinear, pumped generalized bounded-polynomial, pumped max-plus, tracked-primary componentwise pair, tracked-primary lexicographic pair, KO7-specific max-depth, and KO7-specific pure head-precedence. For the pair-valued families, the theorem reads their first component as the tracked primary scalar. The Lean development first proves the scalar-only preliminary ko7_nat_direct_escape_trichotomy; the theorem stated here is the KO7-specific extension ko7_direct_escape_trichotomy_extended, which adds the tracked-primary pair families. A second theorem, ko7_projection_escape_trichotomy, extends the same classification pattern to successful componentwise orienters that admit a fixed scalar projection whose strict decrease is forced by the ambient order and whose projected scalar lies in the listed Nat-valued universe. This is how the weighted scalar-projection and balanced mixed-coordinate matrix families are absorbed into the trichotomy layer. The generalized degree-bounded polynomial family of Theorem 4.6 is now internalized directly into the Nat-valued universe; projection-based matrix families still enter through the separate projection theorem. So the twelve-family barrier package and the twelve-family trichotomy universe are not identical: the former lists the projection-based matrix families directly, while the latter replaces them by KO7-specific max-depth and pure head-precedence and then reintroduces the projection families through the separate representability theorem. Dependency-pair frameworks, generic path-order metatheory, arbitrary semantic methods, and arbitrary mixed matrix orders remain outside this theorem universe.
Table 7: Barrier-package families versus the trichotomy theorem universe.
Shared by both universes Barrier package only Trichotomy universe only
Additive, transparent-compositional, pumped affine, pumped restricted-quadratic, pumped bounded cross-term quadratic, pumped bounded multilinear, pumped generalized bounded-polynomial, pumped max-plus, and the tracked-primary componentwise / lexicographic pair families Projection-based matrix families, reintroduced later through the separate projection theorem KO7-specific max-depth and KO7-specific pure head-precedence
@@
The third disjunct is the explicit twelve-family universe enumeration just listed. Equivalently, every successful direct root-step orienter in the theorem universe escapes by violating wrapper sensitivity on its tracked primary scalar, violating transparency there, or leaving the formalized direct families. Formalized in Meta/EscapeTrichotomy.lean (ko7_direct_escape_trichotomy_extended; scalar-only preliminary: ko7_nat_direct_escape_trichotomy).
Projection-based extension. The same escape pattern also covers the weighted scalar-projection componentwise family and the balanced mixed-coordinate componentwise family. The reason is structural rather than ad hoc: if the ambient componentwise order forces strict decrease of a fixed scalar projection, and that projected scalar already belongs to the explicit Nat-valued direct universe above, then the trichotomy applies to that scalar as well. The Lean proof is ko7_projection_escape_trichotomy in Meta/EscapeTrichotomy.lean, with the concrete representability lemmas matrixFunctional_projection_representable and matrixMix2_sum_projection_representable.
Scope note. The theorem is intentionally restricted to a finite, reviewable direct universe plus a separate projection-based extension principle. This is a deliberate design choice. A theorem over a listed class is verifiable and falsifiable, while an unbounded claim about all imaginable orienters would be either tautological, by defining “outside the formalized classes” to absorb every counterexample, or unprovable with current methods. This is therefore a classification theorem for the listed universe, not a dichotomy between success and failure: the first two disjuncts are testable structural predicates, while the third localizes the remaining open space outside the formalized direct families. The bounded cross-term, bounded multilinear, generalized bounded-polynomial, and max-plus scalar families are included directly; the weighted scalar-projection and balanced mixed-coordinate matrix families are included through the projection theorem; and the current matrix-side strengthening now also exposes fixed-row and row-sum projection corollaries in Meta/MatrixProjectionCoverage.lean. Dependency-pair frameworks, generic path-order metatheory, arbitrary semantic methods, and arbitrary mixed matrix orders remain outside the present theorem universe, even though the paper separately proves negative or positive results for some of them.
@@ 4.8 Scope
Status of full Step. Root-step Step is terminating: mechanized via the nonlinear polynomial interpretation (wf_StepRev_poly) and the KO7-specialized MPO (wf_StepRev_mpo). The full context closure is also strongly normalizing in Lean via wf_StepCtxFullRev_poly, and independently via TTT2/CeTA (§6). Full Step is not locally confluent.
@@
For every trace $a$, the full kernel relation fails local join at $\mathrm{eqW}\;a\;a$: the two one-step reducts $\mathrm{void}$ and $\mathrm{integrate}(\mathrm{merge}\;a\;a)$ are distinct full-step root normal forms.
Proved as not_localJoinStep_eqW_refl in Meta/EqW_Guard_Barrier.lean. The second reduct is in normal form for a simple structural reason: R_int_delta only fires on $\mathrm{integrate}(\delta\,t)$, while $\mathrm{merge}\;a\;a$ is never $\delta$-wrapped, so no rule can act on $\mathrm{integrate}(\mathrm{merge}\;a\;a)$. This is why the SafeStep eqW guards are genuine confluence guards rather than presentation artifacts.
@@
The barrier section isolates a concrete frontier. Inside the formalized direct classes, the duplicating step is excluded by the schema, affine, restricted quadratic, bounded cross-term quadratic, bounded multilinear, generalized bounded polynomial, max-plus, tracked componentwise fixed-dimension vector, tracked lexicographic pair, balanced mixed-coordinate pair, and weighted scalar-projection componentwise barriers, together with the symbolic variable-condition / KBO corollary. The trichotomy theorem classifies the scalar side through max-plus and multilinear, and extends to the weighted-functional and balanced mixed-coordinate matrix families through projected representability, while the generalized polynomial theorem isolates the remaining bounded direct-polynomial escape condition. Crossing the frontier therefore requires one of three moves exhibited in the artifact: projection-style transformation (DP), direct nonlinear escape outside the formalized classes (the global witness $W$ and the KO7-specialized MPO), or ablation of the offending structure itself (removing duplication dissolves the barrier, while a directed wrapper-collapse surrogate does not). The SCC theorems show that this frontier is not confined to immediate single-rule duplication. Delayed duplication across finite cyclic SCCs, and more generally across graph-indexed SCC families once a nonempty closed path, round-trip SCC witness, or bounded finite-round-trip reachability package is supplied, recreates the additive and affine obstruction. Even multiplicity-preserving synchronized cycles still force additive failure, with a conditional affine companion and transparent / scalar-projection downstream corollaries, once the latent packet is exposed under visible wrappers.
@@ 4.9 Orientation catalog
    Path Orders ($\dagger$/L): LPO orients the full TRS (CeTA-certified); this is a framework escape from the $\mathbb{N}$-valued measure classes (Cor. 4.26). Meta/MPO_FullStep.lean proves KO7-specialized MPO orientation (wf_MPORev, wf_StepRev_mpo); this is a fixed-signature comparator, not a generic MPO metatheorem.
Here (L) means Lean-checked inside the artifact, (T) records the corresponding external TTT2 outcome when one is reported, and $\dagger$ marks theory-level or tool-level remarks rather than a new standalone Lean family theorem.
@@ 4.10 Gallery of explicitly excluded methods
Exact polynomial tie (poly_mul_ties_rec_succ).
@@
Formalized as poly_mul_ties_rec_succ. This is not merely non-strict failure but exact cancellation: the duplication of $s$ under app is algebraically absorbed by the multiplicative coupling. Only nonlinear cross-variable coupling (as in $W$) breaks this tie.
@@
The duplication barrier and the flag barrier target different rules and impose contradictory demands on any single $\mathbb{N}$-valued measure: rec_succ requires insensitivity to duplication of $s$ (additive size fails), while merge_void requires sensitivity to flag structure (the $\delta$-flag must strictly decrease). See dual_barrier_rec_succ_and_merge_void.
@@
In a strongly normalizing and confluent calculus with a computable normalizer and decidable equality, the set $\{t\mid t\Rightarrow^{*}c\}$ for any normal-form target $c$ is decidable: compute $\mathrm{normalizeSafe}(t)$ and compare. This is formalized as the Decidable instance reachability_decidable in Meta/Newman_Safe.lean: given $c$ in safe normal form, the predicate $t\mapsto\texttt{SafeStepStar}\;t\;c$ is a DecidablePred, reducing reachability to a single normalization followed by equality comparison. Complexity bounds for related TRS classes are discussed in [3, 7, 16].
@@ 5.1 Strong normalization (SafeStep)
Proof. By exhaustive case analysis over the eight SafeStep constructors, discharged in Lean as measure_decreases_safe_c in Meta/ComputableMeasure.lean. For rec_succ, $\delta$-flag drops ($1\to 0$). For rec_zero and the guarded merge-cancel branch, $\kappa^{M}$ drops in the Dershowitz–Manna order. In the remaining guarded merge / eqW / integrate cases, the first two coordinates tie and the weighted size $\tau$ strictly decreases.
@@
$\texttt{ctxFuel}(\texttt{void})$ $=0,$
$\texttt{ctxFuel}(\delta\,t)$ $=2^{\,\texttt{ctxFuel}(t)+1},$
$\texttt{ctxFuel}(\texttt{integrate}\,t)$ $=\texttt{ctxFuel}(t)+1,$
$\texttt{ctxFuel}(\texttt{merge}(a,b))$ $=\texttt{ctxFuel}(a)+\texttt{ctxFuel}(b)+2,$
$\texttt{ctxFuel}(\texttt{app}(a,b))$ $=\texttt{ctxFuel}(a)+\texttt{ctxFuel}(b)+1,$
@@
$\texttt{ctxFuel}(\texttt{eqW}(a,b))$ $=\texttt{ctxFuel}(a)+\texttt{ctxFuel}(b)+4.$
[…]
@@
For the partial context closure SafeStepCtx, the remaining global local-join obligation is discharged, so SafeStepCtx is confluent. The Newman characterization is retained as a corollary:
@@
Since strong normalization is already proved, this gives both an unconditional contextual confluence theorem and the corresponding SN-plus-local-join characterization for the partial context closure.
See Meta/SafeStepCtx_Confluence.lean (localJoinAll_ctx, confluentSafeCtx, confluentSafeCtx_iff_localJoinAll, and the specialized Newman direction newman_safeCtx).
@@
The Lean proof (wf_StepCtxFullRev_poly in Meta/ContextClosed_SN_Full.lean) lifts the nonlinear polynomial witness $W$ through every constructor position and shows that every contextual contraction still strictly decreases $W$.
@@
Formalized in Meta/SafeStep_Complexity.lean (stepCtxFullPow_length_le_W, stepCtxFullPow_length_lt_W). The result is a concrete complexity theorem for the full contextual relation; the ordinal discussion below remains a separate proof-theoretic calibration for the guarded SafeStep argument.
@@ 5.2 Certified normalizer
Note that normalizeSafe computes normal forms with respect to the guarded relation SafeStep, not the full unguarded Step; a SafeStep-normal form may still admit unguarded reductions. Moreover, SafeStep is deterministic at the root (safeStep_deterministic in Meta/NormalizeSafe_LowerBound.lean), so the certified normalizer follows the unique possible guarded root reduction at each step. Under confluence, reachability reduces to normalization and equality check. For safe normal-form targets, the resulting decision procedure also carries an explicit certified cost envelope: Meta/Reachability_Complexity.lean proves both the linear size upper bound reachabilityDecisionCost_le_linear_termSize and the coarser inherited tower bound reachabilityDecisionCost_le_complexity_bound, together with the matching linear families reachabilityDecision_has_linear_lower_family and reachabilityDecision_has_linear_size_family.
@@
See Meta/NormalizeSafe_LowerBound.lean (mergeVoidChain_pow_void, normalizeSafe_mergeVoidChain, normalizeSafeSteps_mergeVoidChain). This gives a concrete linear lower bound for the cost of the certified root normalizer on one explicit family.
@@
For every term $t$, the deterministic root-step counter normalizeSafeStepstt is realized by an exact guarded root reduction path and satisfies the linear size bound
$\texttt{normalizeSafeSteps}(t)+2\;\leq\;2\,\mathrm{termSize}(t).$
Hence it is also bounded above by the older explicit size-based tower bound inherited from the coarse SafeStepCtx analysis.
The Lean proofs are in Meta/SafeRoot_Complexity.lean; the exact theorem identifiers are listed in the claim-to-code appendix. This yields a fully mechanized algorithmic envelope for the certified root normalizer: an exact linear lower-bound family on merge spines, a linear upper bound in structural term size, and the coarser inherited ctxFuel / cichon corollaries.
In particular, the guarded root relation has a concrete linear lower-bound family for derivation length: each step from $m_{n+1}$ to $m_{n}$ removes exactly one outer mergevoid layer, so the family grows by one constructor layer per step while requiring exactly $n$ guarded root contractions, and $\mathrm{termSize}(m_{n})=2n+1$. Thus the certified root normalizer has linear complexity in the structural size of its input. The coarse ctxFuel and finite fast-growing envelopes remain useful explicit corollaries, and the artifact also exposes the root and guarded-context upper bounds through Cichon-style theorems in the same lex3Note family used by the ordinal calibration.